OpenStudy (ria23):

I need help with this question, it's simple but I can't figure it out and I need it answered quick, help would be much appreciated. Solve the inequality. -1/5x - 4 > - 3 Thank yhu. c:

OpenStudy (ria23):

The choices are between A.) x<35 B.) x>35 C.) x<-5 D.) x>-5

OpenStudy (ria23):

Every time I think I have the answer, I never get one of those, the fraction is confusing me.

OpenStudy (whpalmer4):

\[-\frac{1}{5}x - 4 > -3\]Is that the inequality?

OpenStudy (whpalmer4):

I suggest you multiply everything by 5 to start.

OpenStudy (ria23):

yes.

OpenStudy (whpalmer4):

After you multiply everything by 5, what do you have?

OpenStudy (ria23):

the fraction cancelled and I got, -4>-15?

OpenStudy (whpalmer4):

Not quite: \[-\frac{1}{5}x - 4 > -3\]\[-\frac{1}{5}x*5 - 4*5 > -3*5\]\[-x - 20 > -15\]Right?

OpenStudy (anonymous):

after multiplyin by 5 u will have -x-20>-15 multiplyin again with - x+20<15 notice the sign changes of inequality x<-5 is the answer

OpenStudy (anonymous):

c is the ans

OpenStudy (ria23):

wait, why did the sign flip?

OpenStudy (anonymous):

For an inequality, equal quantities can be added to both sides. SO add 4 to both sides thus giving: \[-(1/5x)-4+4>-3+4\] Thus we get: \[-(1/5x)>1\] Multiplying both sides on an inequality with a negative number leads to a reversal of the inequality. \[(-1).(-1/5x)<(-1).1=-1\] Thus: \[1/5x<-1\] Multiply both sides by 5 retains the inequality: \[(1/x)<-5\] So far so good?

OpenStudy (anonymous):

its the feature of inequality

OpenStudy (anonymous):

by multipling any negative no

OpenStudy (ria23):

OH! I remember that being a step...

OpenStudy (ria23):

RajeshRathod, how did yhu get those steps?

OpenStudy (anonymous):

Which step is not clear?

OpenStudy (ria23):

How did, why did 5 get multiplied by everything? I'm thinking it's because of the fraction... but I don't know... if that's right.

OpenStudy (anonymous):

Just a minute: is it (-1/5)x OR (-1/5x) ??

OpenStudy (ria23):

? There's no parenthesis in the equation.

OpenStudy (ria23):

-1/5x

OpenStudy (anonymous):

Looking at the choices, it has to be (-1/5)x

OpenStudy (anonymous):

So we have: \[(-1/5)x-4>-3\] RULE: Multiplying both sides of an inequality by a negative number reverses the inequality. So in this case multiplying by -1 we get: \[(1/5)x+4<3\] Okay?

OpenStudy (ria23):

Okay.

OpenStudy (anonymous):

RULE: Multiplying both sides of an inequality by a positive number retains the inequality: So, multiply both sides by 5 now so that only x remains. Right?

OpenStudy (ria23):

yea.

OpenStudy (anonymous):

Can you do it now?

OpenStudy (ria23):

yes. Thank yhu everyone. c:

OpenStudy (anonymous):

Signing off - need to go to work. Good luck!!

OpenStudy (ria23):

Thank yhu.